2. Let z = f(x, y) = ln(4x2 + y2) (a) Use a linear approximation of the function z = f(x,y) at (0,1) to estimate f(0.1, 1.2)Please explain b!
Answer #1

passes Eqn of plane = 2x+2y = 2 = 3 Therefore the ear of plane which through the point (a, b, c) will be - 2x+2y-Z = 2 =P whe

2b = 2 fin & 8 a = 2 (ii) 4a²+6² - 4a²+6² Divide both equations : 2b 4a² tb² = 1 = 20 = 1 8a 4a²+6² =) a = b. - putting in (i

Solving both ears give a = 1 & b = 4 which gives c = ln (4x I + 4 = -0.223

Thank you for asking. Please rate my answer so that I can serve you with better answers in the future :) Please comment if you have any doubts.

passes Eqn of plane = 2x+2y = 2 = 3 Therefore the ear of plane which through the point (a, b, c) will be - 2x+2y-Z = 2 =P where a= 2a+2b-c - Since both planes are parallel Now Now, we to the function know this plane is tangent f(x,y) = ln (4x²ty") Now with we take the partial derivation of the f(x,y) respect to a keeping y constant. af (x,y) = OPI - ax ly= b ox ly=b = 2 x8x1 4x² + y² 'y=b 8a = 2 4a²+6² We do the same take keeping & constant of(x, y) dy x=a partial derivative wity = OP I dy 1xza - 2 X=a Lyl ty? 2b ua²+6²

Solving both ears give a = 1 & b = 4 which gives c = ln (4x I + 4 = -0.223